A body of mass 10kg is suspended by inextensible string from a nail 0 and pulled by horizontal force F until the angle of inclination of the string to the vertical is 30° calculate the value of F and the tension T in the string.
1
Expert's answer
2019-05-10T11:24:00-0400



By Newton's law, the body will be at rest if the sum of the projections of forces is equal. From here we get the angle of deviation from the vertical.


"F_T\u200b\u2217sin\u03b1\u2212F\u2217cos\u03b1=0"


Horizontal force balancing load is equal to:


"F=F_T*tg\alpha=mg*tg\alpha=10*9.81*0.577=56.6N"


The tension force of the thread is equal to:


"T=F_T*cos\alpha+F*sin\alpha=85+28.3=113.3N"

Answer: 56.6N, 113.3N